Posts

Showing posts from October, 2025

NPP Projection Close 10/31

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Both evening rounds of  Bugonia in 35mm went smoothly. The print has shed very little but there was a bit of dust throughout and some residue on the gate after each run. This is all to be expected as we continue to show it time and again. The print looks and sounds as great as it did this morning during the tech. For now it is resting on the middle platter with the fabric cover on top. I cleaned the projector and all gears and rollers as thoroughly as possible and sprayed everything down with the compressed air to eliminate any dust from the final show. Breakers and platter are off for the night.  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: I noticed the LampLOC X and Z sensor failure alarms were lit up on the touc...

NPP Projection Admin 10/31

To do this evening/ was completed: 11/7 A1 playlists built. 10/24 content removed from servers. Darkman and Below the Clouds ingested to LMS. New first run letters up on the marquee.  Inspected the last reel of Jaws, back on shipping reel. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Be aware that the racks for Theater 5 have been rearranged and we have to use to access door to get to the projector if we need. There is a 2x4 by this door to prop it open. The racks have been rearranged to be away from the AC pump to avoid water damage in the case of a leak -A Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Close 10/30

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ran some test film off the platter with Alec and Lillian in booth 1...cleaned the projector and the brain/rollers after the film ran out...turned the projector breakers/tree power off...I moved the test film from the middle platter to the bottom one in preparation for tomorrow's Bugonia tech (note: I brought the pucks down and placed them around the print and placed the brain on the top platter) Night of the Living Dead  screening went smoothly...the band made some final remarks before packing up and I put the rest of the gear away back beneath the curtain in theater 1 once they departed (note: some wires may be out of place...I was not sure which cables had been taken from 2 or 4/6 but these are all split between the bins in 1...JBLs are back in 2 along with their power cords and XLRs) Ordering/Receiving (posters & trailers): Additional posters for  ...

Projection Admin 10/30

To do this evening/ was completed: Met with Greg and Jeff from Lincoln Center today to have them in for a while to look at our equipment for sale. They'll likely be taking a lot, even in T2, it's just a matter of breaking things down and figuring out the financial side of it. This is good news! - LH Presentation Issues: Preshow Clips/Playlist: Built the playlist for Shock A Go Go Shorts, which I'll run through as soon as I get another chance. Built the playlist for Kiss Kiss Bang Bang. - LH Added Bugonia masking for T1 and T2, I'll do the rest tomorrow. We'll need to check out the masking for 35mm. Alec and I will both be here tomorrow to make sure. - LH Trailer Updates: Updated volume for PSA_Book_V3 and the Val Kilmer Series trailer (new volumes up on the board in 4/6) plus I reached out to Kat about the Legends of the Fall trailer since it was super low volume with soft dialogue. We now have a V2 version, which also has a different volume to 4.0, so please see t...

NPP Projection Close 10/29

To do this evening/ was completed: Nintendo Switch Private T6 went smoothly. Broke down the HDMI cable and transmitter, those are in 4/6. Disconnected the projector HDMI we used as well. Just a note: we rearranged this connection setup in the booth, disconnected the splitter from everything and just ran with an HDMI from the receiver to the extractor, and the projector HDMI cable connected to the extractor as well.  T2 Halloween drag went smoothly with just a hiccup. The DCP we made for Malice's number featured a slowed-down version of the correct song. I thought maybe the frame-rates could've been mismatched in DCPoMatic but I'm not sure what caused this. So I played the song off the laptop instead, which caused me to be a little imprecise with that number's lighting cues and the transition into the last song as I had to switch back over to digital one while fast-forwarding through Malice's DCP. But overall the show went well.  Ran through T1 platter twice with Ale...

Projection Admin 10/29

To do this evening/ was completed: Staff Screenings are out. Playlists have been built and scheduled on the LMS. Shows are blocked on the calendar.  Presentation Issues: Preshow Clips/Playlist: NOTE - I adjusted the 35mm playlists for Bugonia (which we'll need to do again when the keys open up so we can get the lens/masking right, but I also added 2 minutes of black (with the analogue cue and volume attached) so that those cues aren't attached to the DCP. The DCP should be ever so slightly behind the print, too, in case something goes awry and we need to switch over to DCP.  Trailer Updates: Special Events/ AV set up: External speakers, Yamaha, Wireless mic receiver and such are setup in T2 for the drag show tonight.  For the private event (Nintendo Switch situation) in T6 tonight, I brought up the hdmi transmitter and receiver from T3, using the splitter in 6 and connected everything in the booth and at the stage for audio and image. I built a playlist and the cue for ...

NPP Projection Close 10/28

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Scheduled regular SPLs for Depeche Mode: M on 10/30 and 10/31 in theater 3 and adjusted the cue sheets accordingly. Double checked this and adjusted the credits light cue as there was a cleaning cue at the top of the credits in place of the normal pre-show level one. Added generic welcome/feature/policy stingers prior to the feature so the feature will start thirty minutes in. This playlist does not have any trailers Trailer Updates: Special Events/ AV set up: Raised the volume to the feature to compensate for the bigger crowd but otherwise the  Depeche Mode: M screening went smoothly. Intro and Q&A went well. Mics, chairs, and stanchions were returned to storage following the show. Stanchions are back behind the curtain Set up for trivia at Trees Innocent Blood went very well, I rewound the print heads out on house reels for reinspection since it's an Academy print. -A Ordering/Receiving (posters ...

NPP Projection Admin 10/28

To do this evening/ was completed: Kiss Kiss Bang Bang and Prince of Egypt features ingested, playlists built for Prince of Egypt, Nashville, Color Purple, Sundays on Fire, Nouvelle Vague. Looked at Halloween and Depeche Mode, lights up cues set. Color Purple and Kiki's slides made.  Bugonia print arrived, resting in T1. Color Purple on house reels. Presentation Issues: Shelby Oaks OCAP playlist had the flat DCP with a scope lens/mask pack, patron notified MOD some subs were getting cut off, quickly switched to flat lens and mask about four minutes into the show, updated the playlist.  One Flew Over the Cuckoo's Nest DCP downloading via Aspera, could finish tonight, keep an eye. Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ran loops for tonight's show and focused and masked the lenses. - MC Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Returned the flash drive with Imaginationland on it back t...

NPP Projection Close 10/27

To do this evening/ was completed: Spook Show performance went smoothly. Equipment broken down afterward.  Built a 35mm Bugonia playlist that's just a dowser close cue at the end of the preshow and a cue to switch to multi-channel analog. Scheduled the week in the LMS with some exceptions. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 10/27

To do this evening/ was completed: See AV notes below!  Presentation Issues: Preshow Clips/Playlist: Cristina and I tech'd the rest of the Shock A Go Go shorts. I will build out this playlist. - L Adjusted the 10/28 Depeche playlist for intro/q&a and made a 10/30 (which can also be used for 10/29) playlist without any intro/q&a - L Trailer Updates: Special Events/ AV set up: As much as I would have liked to do a run through in advance with the Spook folks, they're from out of town, so that was kind of out of the question anyway, but we made it work. The lav mic I tested in T1 was having issues with static and feedback in T2 no matter how much I adjusted it. We wound up getting another from one of the talent members and it's working well! As backup, we will have two wireless mics (one of which we already have for the show anyway) and a mic stand off to the side. We went through the whole show (they forgot to send the intro sound cue until we were teching, so I DCP it...

NPP Projection Close 10/26

To do this evening/ was completed: Inspected the rest of Nashville and started on Jaws, good condition. Transferred all of the Spookshow preshow pieces to T2, all of the audio file DCPs start with Z (Z1, Z2 etc). Reached out to Universal/DHL for Casper pickup, projection CC'd. New 35mm letters up on the marquee. Looks like no lights up cue for Jurassic Park   III if someone can add that tomorrow. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Open 10/26

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Tech'd Creature from the Black Lagoon and Indiana Jones 3  in their respective theaters (connected the mic in 3 and brought the spotlight remote down from 4/6) and their was no issue with the content...naturally the hosts in 3 opted to not use the mic and spoke to their guests without it Ran loops in theater 1 prior to the Casper screening which went smoothly...placed the print back on its shipping reels and zip tied everything shut...this is at the bottom of the steps in booth 1 Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Deleted  Creature from the Black Lagoon and Indiana Jones 3  features from the servers after the shows concluded Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Two bulbs were out in theater 3 this morning...one needed replacing and the o...

Projection Close 10/25

To do this evening/ was completed: Inspected all 9 reels of  The Color Purple , which is structurally a little on the rough side, plus the cues might be a little challenging, but it should run just fine.  Presentation Issues: No issues this evening.  Preshow Clips/Playlist: Built the playlists for Depeche Mode and Night of the Living Dead. Checked the others.  Trailer Updates: Special Events/ AV set up: Friendly reminder we've got the private of Indiana Jones tomorrow morning! There's an intro. Mics in T3 already.  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: 64 has been frigid in the booth office, so I brought the temp up.  Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Report 10/25 Open

To do this evening/ was completed: Presentation Issues: There was a douser close cue placed right before the Creature from the Black Lagoon feature in the playlist, despite there not being any lens/masking change. I was radioed once the movie started that the screen went black. Quickly checked the VNC and noticed the douser was closed. I opened it, of course. I think it was early enough into the movie that we didn't need to rewind. I updated the playlist for tomorrow. Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: The private event in T3 went well, no issues of any kind. Ditto the intro & screening of Casper in T1, it was as smooth as could be. The print is rewound and ready for tomorrow and the projectors are clean. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested the NOTLD DCP to the LMS. Heads up, I just renamed it ("NOTLD-LSC_FTR...") so it's clear it's for Live Sound Cinema as we'll want to ke...

NPP Projection Close 10/24

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Began transcoding Night of the Living Dead in DCP-o-matic for its live music screening on 10/30 Outgoing/Shipping (content): Print/video Inspection: Continued inspecting Nashville this evening Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Admin 10/23

To do this evening/ was completed: Just fyi - before anyone has questions about plattering Bugonia and our regular projection schedule, Alec and Matthew and I are working on this to see what we need overlap for and will keep you abreast of any changes, but for now, it's really looking like it just be me shifting to a mid for that opening week. A lot of this depends on how well it does and how much we'd like to schedule, but I'd also like to honor the fact that y'all have stuff outside of work so we'll make sure we're covered and happy.  - Lillian A1 10/31 playlists built-- Bugonia does not yet have the proper masking/lens presets included heads up. 10/17 playlists removed from servers. New posters up, outgoing posters down. Nashville and Jaws prints arrived, in T1. Inspected the first two reels of Nashville, decent condition but lots of splices. Playlists made for Indiana Jones, Bride of Frankenstein (incomplete, will need updates). Made DCP of all the performa...

NPP Projection Close 10/23

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Built a show playlist for Innocent Blood  (including its custom intro slide) but the volume cue change for the feature will need to be added Trailer Updates: Special Events/ AV set up: Brought up a mic stand to theater 6 for the private event intro + the feature needed a volume bump but overall the  Return of the Living Dead screening went smoothly KGO people did a quick tech of their own since everyone who was absent this morning was available to breeze through the presentation...this ran two minutes over schedule so I had to start the show playlist manually...followed along with the ROS sheet to make the volume/lights cues as accurate as possible. The rest of the screening went smoothly. AJA and mic equipment was broken down and stored neatly beneath the curtain Ordering/Receiving (posters & trailers): Blue Moon , Pillion , and Shelby Oaks posters were delivered and stored behind 3/7...I lef...

Projection Admin 10/23

To do this evening/ was completed: A lot of teching and a lot of meetin' and greetin'! Whew. See all notes below, but for tonight (when there's time!) please start building more SE playlists. That would be enormously helpful! Thank you! Can't wait to get rid of some more stuff on the server tomorrow... - Lillian Presentation Issues: Preshow Clips/Playlist: Ran through all the new playlists and their content - Frankenstein (I slightly adjusted volume in one preshow segment for this), Springsteen all techd, Shelby Oaks and Blue Moon sorted as well. Now, I did notice that the A1 playlists had the scope version of the trailer for Sentimental Value (please be careful when ingesting/building!) so I switched it to Flat. - L Shleby Oaks feature should be in Scope. Playlists updated.  L Frankenstein feature is not Scope, and now is in Flat. Updated. - L Trailer Updates: Special Events/ AV set up: Finished the setup for KGO tonight - mics are charging. Added a new spots setting f...

NPP Projection Close 10/22

To do this evening/ was completed: Inspected Innocent Blood-- excellent condition, on house reels. Scooby Doo went smoothly, packaged it up afterward, mics returned to the booth. Warriors print on the side entrance cart. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 10/22

To do this evening/ was completed: Screenwriter rebooted this morning - Christie contact made for T1. No issues.  Had our cinema meeting today and we're doing extremely well on assets for this week. Spook show needs are going well. Sorting out the details with MOD.  I came in this morning to find many of our servers with very little space, so I did a big write up of what needs to stay and what could go depending on the screen/showtimes.  We're in the clear now!  Presentation Issues: Preshow Clips/Playlist: Staff Screenings have their playlists built, scheduled, blocked on the calendar.  Trailer Updates: Special Events/ AV set up: Partially set up for KGO tomorrow with the rest of the equipment ready to go (AJA, etc). Wireless mic mixer and Yamaha are sorted, I mainly wanted to test out the LAV mic as soon as possible (for Spook show). Granted, the spook show will be in T2 and not in T1, but still. Changed the batteries in both the receiver and transmitter and th...

NPP Projection Close 10/21

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Cristina dropped off a case of lavalier mics for Monday's tech/magic show Touched base with the hosts of The Deuce and went over the run of show for this evening...the pre-show ran 20 minutes behind schedule since the intro/slideshow took way longer to finish despite the hosts telling me it wouldn't last more than ten minutes which, followed by the raffle, pushed the film's start back longer than normal. The screening itself went just fine. Most of the audience left as soon as the credits started so the Q&A happened, albeit sadly. I cleaned the projectors once they'd cooled down, brought the mics back up, and put back all the chairs/neatly stored all the gear used back beneath the curtain. The print is tails out in booth 1 and can be rewound tomorrow. Ordering/Receiving (posters & trailers): Pre-Code Parade authors' book sale standees w...

NPP Projection Close 10/20 + Admin 10/21

To do this evening/ was completed: Finished inspecting Dr Moreau, good condition, on house reels 4/6. Ingested the Imagination Land slide and ran the new DCP for the filmmakers, they're happy. Ingested Blue Moon and Shelby Oaks--Shelby came with two separate DCPs, one flat, one scope. Apparently there are ratio changes in the film so they provided two versions to accommodate theaters who don't alternate between ratios. Assuming we'll go flat so that's the one I built the playlists with. Set up the mic and image connections in T1 for tonight's Warriors show--all is working. Muted sound on the Yamaha mixer so just turn that volume back up when the time comes. Aux cord is set up in Trees for the DJ set before the Deuce. Jeff just goes off his laptop and it should be fairly simple. The mixer under the bench is already flipped on.  Scheduled the coming week in LMS with a few exceptions-- Depeche Mode needs a playlist made. Tech'd all of the Halloween dragshow pieces...

Projection Admin 10/20

To do this evening/ was completed: For tomorrow (and see notes below in SE, too!) Friendly reminder that there are two techs on the calendar, one of which is for Imaginationland (private event this weekend) in t3. You can just get the DCP setup for them to watch (masking/lights/volume, etc) and leave them with it until the very end. Please use the DCP with 1017 in the CPL. They didn't provide a new file with a V2 label, so look for the file with 1017 in it! This will cross over into tech with Cristina for 35mm tomorrow, but you guys (Jack and McKegg) can each do one, plus you don't need to stay with the Imaginationland folks tomorrow. Chenelle will be there FYI. She's sent out a calendar invite.  Presentation Issues: Preshow Clips/Playlist: Looked at Hellrasier and Demon City (both of which look and sound good). I raised the volume considerably for Hellrasier , now at 7.0.  Deleted old playlists and content. Please keep building what can get built and we should be receiv...

NPP Projection Close 10/19

To do this evening/ was completed: Inspected Night Nurse-- good condition, back on shipping reels for now. Brunch letters up on the marquee. Judi dropped off the thumb drive, ingested the new DCP of her film to LMS and T3. Hellraiser finished downloading--I ingested it to LMS and then to T1, which took a while, just finished at the end of the night so will have to tech tomorrow. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Open 10/19

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Beetlejuice  and I Still Know What You Did Last Summer played in theaters 1 and 4 without issue (removed these features from their respective servers after the shows concluded) Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Began downloading the Hellraiser  DCP file via MASV to the 4/6 desktop Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Air conditioning units in booths 3/7 and 5 were down this morning...alerted the MODs and left the doors to these spaces open to let the hot air out Cue sheets updated/checked (every shift): Marquee updated?: Removed the NewFest letters from the brunch side and organized the letter cabinet behind 3/7 Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Close 10/19

To do this evening/ was completed: Swept the LMS of NewFest content! Woohoo! Great job, y'all. Our team really rocked it this year and I am very pleased. Thank you for helping to run such great shows. NF folks are really happy.  Presentation Issues: Preshow Clips/Playlist: Deleted old playlists and got rid of New Fest shorts, features, slides, you name it.  Got my eyes on I Still Know What You Did Last Summer . Playlist is all set! Trailer Updates: Special Events/ AV set up: We are still waiting on the DCP of Hellrasier. Even before I came in I was emailing with the guy who's getting this DCP to us, which failed on him once and he is, apparently, working remotely from a film festival. I asked if he could send it to us using a different transfer service and provided a few suggestions. I wrote back to him asking for an update as he mentioned MASV and wetransfer (and wetransfer can be spotty for large files...) No word yet. Depending on what we hear by the morning, I'll make ...

NPP Projection Open 10/18

To do this evening/ was completed: Looked at Josie before NewFest--both events went smoothly, mics back where they belong. Ditto Trees event. Looked at Beetlejuice in between rounds--volume and lights up cue all good. Downloaded/converted/ingested all of the musical numbers for Miss Malice Halloween--curtain call number has the tips slide attached. NewFest posters taken down, replaced with some Coming Soons. Also put up a  Bugonia   poster outside. Per Matthew's request, No Kings up on the 35mm side of the marquee. Return of the Living Dead still transcoding in T1 but should be done very soon. Continued downloading SAGG shorts-- Feast Serpent and Something Unforgivable ingested to LMS. Sasquatch Room downloading to Share drive. Handed Judi's thumb drive back over. She said she'll back tomorrow or Monday with new content to ingest. Built some SEs--both Caspers and Creature from the Black Lagoon. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV...

NPP Projection Close 10/17

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Built and scheduled a show playlist for Demon City Shinjuku on 10/20 (this is set to transfer to theater 4 on 10/19 + still needs credits/volume checked) Built all first run show playlists for the week of 10/24 with the exception of Blue Moon and Depeche Mode: M  (waiting to receive feature content for these titles) Double checked NewFest playlists for tomorrow afternoon's feature and shorts programming and everything is in order Trailer Updates: Special Events/ AV set up: Placed the turntables back in the case along with the mixer and sealed it shut/stored it neatly beneath the mezzanine steps Ordering/Receiving (posters & trailers): Die My Love , Hamnet , and Jay Kelly posters were delivered and stored behind 3/7 Ordering/Receiving (content): Prints of Night Nurse and The Color Purple  were delivered and brought up to booth 1 Currently no sign of the Hellraiser  feature file but will con...

Projection Admin 10/17

To do this evening/ was completed: ALMOST done with NewFest - let's not forget to be on the lookout for other special events we have coming up. October is a busy month for us, so please be sure to check the calendar, your schedule, and make note of the shows you'll be working.  Presentation Issues: Preshow Clips/Playlist: Tech'd and updated the NewFest playlists for this week with the exception of the DCP of Josie and Pussycats. The KDM has not opened yet. Please see info below in Ordering/Receiving. Added house lights info for Josie and Beetlejuice but these DCPs still need to be tech'd! KDM info below.  Trailer Updates: Special Events/ AV set up: Finally got looped into the thread for Hellrasier and made contact with the distro folks who are re-sending the DCP as the transfer linked failed (on their end) yesterday. This should show up today. I will keep y'all abreast of any new info. I asked if they need our certs for the KDM. Update: DCP will be unencrypted. Fea...

NPP Projection Close 10/16

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Tested spots and set up mics (only one of two was wired) in theater 3 between rounds prior to the NewFest feature pre-show starting  Brought the wireless mics down from booth 1 and tested these before the shorts program began + placed one on a stand for the intro and left the remaining three on a table for the filmmakers to use for the Q&A Set up the mixer at Trees Lounge with Lillian's help + last minute request from the NewFest people for a microphone by the step and repeat which I wired and placed next to the red carpet Mixer event at Trees went smoothly...DJ stuff was put back under the mezzanine steps (NewFest gear is still there and in the way so the turntable case can get covered and properly stored some time tomorrow) Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Insp...

Projection Admin 10/16

To do this evening/ was completed: Cristina sent the folder for the Shock A Go Go shorts (which have an upload deadline of 10/20 for the filmmakers). There are already two shorts in there from what I could see, so I made a folder on the desktop (preshow/trailers drive) to download the files and DCP what needs to be DCP'd. I have the first one, Parasomnia, currently transcoding in 4/6. Bloody Tears mov file is downloading to the desktop folder. For the DCP container name, please do this: (Title)_SAGG. I think it'll help to organize, so that we know it's a Shock A Go Go short. Thanks! - Lillian  More SE playlists built and scheduled: Beetlejuice, I Still Know What You Did, Hellraiser (just needs feature added, Scooby Doo. Inspected reels one through four of Casper, good condition overall, some decent wear on reels three and four, emulsion scratch at the end of reel four. On house reels. Just one reel left to inspect. Presentation Issues: Preshow Clips/Playlist: Looked at Afte...

NPP Projection Close 10/15

To do this evening/ was completed: Both NewFest events went smoothly this evening. Mics have been put away. Built the coming week's cue sheets.  Jeremy W came by to pick up his DCPs, handed them off in the lobby. I Still Know What You Did... ingested to LMS. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 10/15

To do this evening/ was completed: There is a backup file for Lesbian Space Princess (one of the NF shows tonight) but there is no need for it. I have forwarded the info to Jack, but I ran through the DCP this morning for the fourth time. This is just in case of emergency.  Presentation Issues: Preshow Clips/Playlist: Updated the If I Had Legs playlists with the Innocent Blood trailer (as per Kat's request this afternoon) once the Hellrasier playlist can get built, we'll make sure the trailer is in there too.  Trailer Updates: Special Events/ AV set up: Tech'd both NewFest features today and checked in with the coordinators to confirm everything. Also sent Rachel images of the Audience Award slides for All About the T and Ninxs to let her know that theyr'e correct (for tomorrow). Mics are all set in both theatres. Two wired in T3, wireless in T1. If you need to abandon the PC in T1 because the wifi has gone off, you can always hook up the rca to the iMac! For now it...

NPP Projection Close 10/14

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: The feature screening in theater 3 went smoothly but I never got word from any of the NewFest organizers that it was ending or had ended until after I left the booth around 830PM to see about the DJ setting up at Trees Lounge. Thankfully there was time between rounds so I was able to place the mics and stands back beneath the curtain in the time it took for the final short in theater 1 to conclude.  The DJ did not need any assistance setting up but did make a comment about the wires showing (???) and being unsightly which I couldn't do anything about since the table cloth was already being used to cover the DJ case/table contraption and to cover the wires would mean covering the mixer/CDJs (priorities!!!) As expected, the NewFest folks had the bright idea of wanting to tech the show playlist for the New Voices shorts program between rounds (which was in direct...

NPP Projection Report 10/14 Admin

To do this evening/ was completed: Scheduled the coming week in the LMS. There are a few SE exemptions, playlists still need to be made. Inspected reels two through five of Blade of Fury-- all reels feature English subtitles. There was a small section or two where I saw Chinese characters with no English beneath them on reel three. Print is on house reels. Tech'd both of tonight's NewFest programs with Desmond, good to go. Presentation Issues: Preshow Clips/Playlist: Per a request from NewFest, I increased the volume on all forthcoming shorts +1.0. I also raised the volume of their trailer to 4.5 for T3 and 5.0 for Theater 1 in all playlists. Trailer Updates: Special Events/ AV set up: I set up the DJ equipment for the party in Trees tonight. The CDJs are connected to the mixer and that's connected to our sound system. I tested everything and it's good to go. There's also a 1/8" cable if they want to connect their laptop as well. Plus an ethernet hookup should ...

NPP Projection Close 10/13

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Built and scheduled a playlist for the  Frankenstein tech tomorrow morning (no pre-show just a feature transitional) in theater 4 Double checked content and volume cues and scheduled both NewFest playlists in theaters 1 and 3 for tomorrow evening (New Voices short program + Serpent's Skin feature film) Trailer Updates: Special Events/ AV set up: Set up a mic for NewFest next to its step and repeat stand at Trees Lounge + I left this plugged in but turned off the mixer since there is a mixer tomorrow night and chances are they'll ask for it to be set back up again (DJ key is back in the A/V bin in the mezzanine closet) Same Again intro and screening in theater 6 went smoothly (I told the organizers we would not have time to run through any playlist or feature content given the short flip but did tech the mic and spotlights before the show began) Here Come the Dolls shorts program ran fine save for th...

NPP Projection Report 10/13 Open

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Will is right, we should be on Non-Sync in all theaters for playing Spotify in the house. Digital 4 is for sound from HDMI connections and only in Theaters 1 & 2. I noticed this on Saturday, updated the playlists for that day, then forgot to mention it after everything wrapped up. I've adjusted the cues in all NewFest playlists. Tested in both 1 and 6, all set. We got a note from Chrissy at NewFest about the aspect ratio of She's the He opening at the end credits. I checked it and added a cue to open the masking for the credits to 1.78. I also set it to play entirely in the 1.85 lens file so nothing gets cut off. Also ran through both NewFest playlists for tonight, made some masking/lens adjustments to a few of the shorts in 1. Otherwise everything was in order. Desmond suggested we bump the volumes for the New Voices shorts (playing tomorrow night in T1) as Monisha from the fest will be here a...

NPP Projection Open 10/12

To do this evening/ was completed: Presentation Issues: I noticed during my techs this morning that the NewFest show playlists are set to start on the "Digital 4" channel for the pre-show Spotify playlist, however, I only got playback once I switched to the "Non-Sync" channel...future playlists could benefit from being updated with this cue so as to keep us from having to switch channels manually at the top of every program (I've added this cue to the shorts program and feature playlists for tomorrow evening but please double check this is working tomorrow morning) Preshow Clips/Playlist:  Trailer Updates: Special Events/ AV set up: Set up and tested mics in theaters 2 and 3 prior to the events (brought an additional mic stand down to theater 3 for the intro/Q&A and took both up to the 4/6 booth after the event + returned all mics to their charging ports) In Your Face shorts block and Summer's Camera screenings went smoothly (note: keep the audience awar...

Projection Close 10/12

To do this evening/ was completed: NewFest is on! We're off to an awesome start. I double checked the playlists for tomorrow. Please run through everything tomorrow, first thing in the morning, then I'm sure there will be another tech beforehand so that everyone's good. Reminder that I put the centre spots info in the event memos that need it. Be sure to test Spotify in the morning for posterity.  Presentation Issues: Preshow Clips/Playlist: Friendly reminder that I checked My Beautiful Launderette and built the playlist for this week. The lens/masking cue for 1.66 is sorted for T3.  Built the playlist for She's the He Also built the playlist for Josie and the Pussycats , but we still need the keys for it. Additionally, there's no mention of the NF logo slide at the end of the feature, so I've left it without. We'll double check with them before the date. The keys come in on the 17th!  Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters ...

NPP Projection Report 10/11 Open

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Both NewFest screenings went very well today! Heads up, they'll need a mic set up in Trees for most if not all of the days they're having screenings here. I left a stand in the storage closet. I re-exported the QMM slides because they were actually not the right size, they should look stretched on the computer so they look right when squeezed by the tube TVs in Trees. The T4T one I made was the right size (weird thing with the video software I was using where it squeezed the slides back to their original size after I stretched them with HandBrake, but the software allows you to resize which I figured out for T4T on Thursday night but I guess it didn't occur to me I'd have to redo QMM as well). I added the stretched T4T and QMM slideshow loops to Plex so we're good to go for this! Ordering/Receiving (posters & trailers): Ordering/Receiving (...

NPP Projection Close 10/10

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Built all first run show playlists for the week of 10/17 with the exception of If I Had Legs I'd Kick You as we are still waiting on pre-show content for it from Kat  Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): New Avatar: Fire and Ash  and  Zootopia 2 posters were delivered and stored behind 3/7 Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Yes Projection cart by entrance?: Yes Additional Notes: Added new first runs to the dry erase board in 4/6

Projection Admin 10/10

To do this evening/ was completed: Matthew flipped the breaker back on for T1 (which is in 4/6 server station) for the PA speakers. I tested mics multiple times and they're definitely working and, more notably, rather  hot . For intros/Q&As, we can bring down the usual level for the mics (on the cp750) from 7.0 to 6.5 or 6.0, I reckon. I did not have the time to get the massive ladder out to adjust the actual speakers, but I think it's worth doing, just so everything's a little even keel.  I added the the Queer Media Mixer slideshow that Alec made to a folder on Plex (which looks great! I tested it!), but the T4T one (for some reason) did not export with the right size, so I'm exporting another one right now to the desktop named T4T_Mixer_133_5hrs. Not sure if it's actually 5hrs but it is certainly more than 3. In any case, it should (hopefully) be the right size. Should be done by late tonight.  Presentation Issues: Preshow Clips/Playlist: FYI - NF preroll loop...

NPP Projection Report 10/9 Close

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Spoons set up was a little hectic as the outlet which the speakers are plugged into was not giving any power. I briefly tried looking for the breaker to see if it had tripped but in the interest of time I just grabbed a power strip and extension cords and connected the speakers (plus the string lights for their tables) to the outlet by the masking motor. I looked around for the breaker after getting everything settled but did not see anything that looked tripped. I'll ask Matthew where the breaker for this is and, if it's not a tripped breaker, have Vlad come look at it. *Also, they like having the Yamaha mixer on the table with the spotlight controller and laptops. Can we make sure this is set with everything else during the day? Other than that, everything was in place which helped a lot with the tight turnaround. **One more thing: Michael mentioned the ...

Projection Admin 10/09

To do this evening/ was completed: Hey ya'll - just a friendly reminder that you gotta ingest the OV before a version file (VF). Frankenstein has been ingested completely.  Presentation Issues: Preshow Clips/Playlist: Techd all new A1 playlists.  Adjusted volume for the Casper trailer and 5min preshow Deadly Games (was a little loud).  Trailer Updates: Special Events/ AV set up:Yamaha, Wireless Mic Mixer, Aja, Spots Board all setup in T1. Aja with hdmi and spots board are resting on the table to the right with three mic stands ready to go.  Confirmed with Galen that the Gilead trailer we techd the other day is OCAP.  As for NF tomorrow, I'll be in early tomorrow to go through the All About The T playlist (double checking that the right VFs are in there - ignore the double of Purebread in there, that's intentional) as well as the New Voices playlist. I thought we had the ratios listed for the preshow/slides (most of which should be 1.78), so I'm also going to dou...

NPP Projection Close 10/8

To do this evening/ was completed: 10/10 playlists updated with 3.7 volume cues (welcome and feature).  Supernatural on the cart with the right shipping label attached. Beautiful Laundrette ingested to LMS and T3. Frankenstein ingested to LMS. New 35mm letters up on the marquee. Bones went smoothly. Mic and cable broken down afterward, mic stand in 4/6. Print has been put back on its shipping reels. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Report 10/8 Admin

To do this evening/ was completed:We're probably getting Bugonia on 35mm. I'm giving the platter in T1 some tests. So far so good. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ran loops for Bones and set up a mic for the intro. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested If I Had Legs I'd Kick You , After the Hunt , and regular 7.1 versions of House of Dynamite . The Atmos version was ingested for the latter. This would play fine, but since we're not an Atmos theater we should just ingest the regular 7.1 version. Ingested Josie and the Pussycats . I started transcoding My Beautiful Laundrette . Since this is a file we don't need to use MakeMKV, that's only for ripping blu-rays. I trimmed out the logos from the beginning with DCP-o-matic (a Criterion logo and a newly added MGM one) so we don't need to have the dowser closed at the start. Outgoing/Shipping (content): I delete...

NPP Projection Close 10/7

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Set up for trivia at Trees Lounge this afternoon Supernatural intro and screening went smoothly, save for some fussing over the volume during the show as the mono soundtrack was either too low at its QC'd volume or too loud at anything above 10 on the 500 and 750. Lillian was texting about level changes during the show. I adjusted accordingly but the volume dial on the 750 is super finicky and will drop the volume way too low or way too high depending on how much force there is applied to it. I think moving forward we should simply split the difference on volume during techs considering how much these print screenings fill up (this one was part of the Pre-Code Parade series) Ordering/Receiving (posters & trailers): Additional Frankenstein  character posters were delivered and stored behind 3/7 Ordering/Receiving (content): Began running the Indiana Jones a...